I would really appericiate if you wouls help me on this problem...
Two thirds of a number and three is the same as 3 multiplied by 1 less than twice the number. What is the number?

Thank you very much

Answer by Nate(3500)

Define the number as: x
2x/3 + 3 = 3(2x - 1)
2x/3 + 3 = 6x - 3
2x + 9 = 18x - 9
18 = 16x
9/8 = x